OpenStudy (anonymous):

Part of $3000 is invested at 12%, another part at 13%, and the remainder at 14%. The total yearly income from the three investments is $400. The sum of the amounts invested at 12% and 13% equals the amount invested at 14%. Determine how much is invested at each rate.

OpenStudy (snowsurf):

Lets start picking some variables here. We have three unknowns. Let use x , y and z. Let x = amount invested at 12% Let y = amount invested at 13% Let z = amount invested at 14% We know the total amount invested is x + y + z = 3000 Each investment took a portion of that money so they sum up to 3000. And the other equation is x + y = z. This comes from the fact that we were told that 12% and 13% invested equals the sum of 14% invested. We know the total amount invested is 400 and that comes from all three investments. 0.12x + 0.13y + 0.14z = 400 So now you have three equations that can be used to find three unknowns.

OpenStudy (snowsurf):

So lets label equation a.) x + y + z = 3000 b.) x + y = z c.) 0.12x + 0.13y + 0.14z = 400 If you take equation b and plug it into equation a you get the following 2z = 3000 Now we can find what z is. Solve for z z = 1500 Now we need to find x and y. Since we know z we have equation a as x + y + 1500 = 3000 x + y = 1500 Solve for y ---->>>> y = 1500 - x Now plug this into equation c.) 0.12x + 0.13 (1500 -x) + 0.14 (1500) = 400 0.12x + 195 - 0.13x + 210 = 400 - 0.01x + 405 = 400 - 0.01x = -5 x = 500 Now we know both x and z. x = 500 and z = 1500 So find lets find y. Let's go back to equation a.) which is x + y + z = 3000. Plug in x and z and solve for y. 500 + y + 1500 = 3000 y + 2000 = 3000 y = 1000 So we have all three variables found. x = 500, y = 1000 and z = 1500.

OpenStudy (anonymous):

Why is this bumped? Wasn't is already solved?